OpenStudy (aravindg):

33. For what value of x, is the value of (x) = x2-4x + 3 is 0?

OpenStudy (anonymous):

the values are simply 1 and 3

OpenStudy (aravindg):

arent they the roots???

OpenStudy (anonymous):

yes they are the q is to find the roots indeed

OpenStudy (aravindg):

thx guys

OpenStudy (angela210793):

D=b^2-4ac=16-12=4 x1=(-b+sqrD)/2a=4+2/2=3 x2=(-b-sqrD)/2a=4-2/2=1